OpenStudy (anonymous):

Is my answer correct?? Find the area of half the circle given the following: Round your answer to two decimal places. Use 3.14 for pi. d = diameter d=16 my answer rounded to two decimal places : 100.48

OpenStudy (anonymous):

YES! :))))

Directrix (directrix):

Area of half circle: 1/2 pi r ^ 2 = 1/2 pi (8)^2 = 32 pi = 100.53 square units (approx)